OpenStudy (anonymous):

Fool's problem of the day, what is the remainder when we divide \( \large 9^{{100}^{{101}^{102}}}-5^{81^{82^{83}}} \) by \(13\).

OpenStudy (asnaseer):

Initial thoughts: 1. \(5^x\) will always end with the digits 25 for \(x\gt1\) 2. \(9^x\) ends in the following set of cyclicly recurring digits:\[\begin{align} 9^1&\to01\\ 9^2&\to81\\ 9^3&\to29\\ 9^4&\to61\\ 9^5&\to49\\ 9^6&\to41\\ 9^7&\to69\\ 9^8&\to21\\ 9^9&\to89\\ 9^{10}&\to01\hspace{2cm}\text{(pattern starts repeating)} \end{align}\] Now in \(100^{101^{102}}\), will result in a number that is evenly divisible by 10, therefore \(9^{100^{101^{102}}}\) must end in the digits 01. Not sure of the next step yet but I hope I am going along the right lines.

OpenStudy (asnaseer):

Another thought on this - using the binomial expansion:\[(a-b)^n=a^n-C_1^na^{n-1}b+...+(-1)^nb^n\] then we can write:\[9=13-4\]and\[5=13-8\]so:\[9^n=(13-4)^n\]\[5^m=(13-8)^m\]therefore:\[9^n-5^m=(13-4)^n-(13-8)^m\]and if these are expanded using the binomial expansion, then all terms except the last one in \((13-4)^n\) will be divisible by 13 and, similarly, all terms except the last one in\((13-8)^m\) will be divisible by 13. We can therefore just look at the last terms in order to work out the remainder when dividing by 13. Therefore:\[9^n-5^m\text{ mod }13=((-1)^n4^n-(-1)^m8^m)\text{ mod }13\]\[\qquad=((-1)^n(2^2)^n-(-1)^m(2^3)^m)\text{ mod }13\]In this case \(n=100^{101^{102}}\) which is even and \(m=81^{82^{83}}\) which is odd, therefore we get:\[9^{100^{101^{102}}}-5^{81^{82^{83}}}\text{ mod }13=((2^2)^{100^{101^{102}}}+(2^3)^{81^{82^{83}}})\text{ mod }13\] Need to think if this is a fruitful avenue to pursue...

OpenStudy (anonymous):

Keep trying, Asnaseer you will get it :)

OpenStudy (anonymous):

Let\[9^{100^{101^{102}}} = b_{1} \mod 13\]\[5^{81^{82^{83}}} = b_{2} \mod 13\]Therefore,\[9^{100^{101^{102}}} - 5^{81^{82^{83}}} = b_{1} - b_{2} \mod 13\]With that established we could get the individual modulus then subtract them. Note that\[\forall a_{1}a_{2}b_{1}b_2 \in \mathbb{Z}(a_{1}a_{2} = b_{1}b_2 \mod n)\] and \[9^{100^{101^{102}}} = 9^{100 * 101 * 102} = 9^{1 030 200}= 9^{85 850 * 12} = 1 \mod 13\]\[5^{81^{82^{83}}} = 5^{81 * 82 * 83} = 5^{551 286} = 5^{49940 * 12}5^{6} = 5^{6} \mod 13\]by repeated use of Fermat's Little Theorem and the property above. Therefore,\[9^{100^{101^{102}}} - 5^{81^{82^{83}}} = 1 - 5^{6} \mod 13 = -15624 \mod 13 = 2 \mod 13\] Therefore, the remainder is 2.